16x^2-4x+4=0
What is the alue of the discriminant?

Answer by xachu4u(10)

Ans:
D=b^2-4ac.
Here b=(-4), a=16, c=4
=> D = 16 - 4*16*4
= 16 - 256
= -240
Since D<0 the roots are imaginary